Use the following information to answer Question 25 to Question 29: The ages of 9 college students who are enrolled in statis28. (1 point) IQR of the data is th 29. (3 points) Identify all the outliers (if any).
Answer #1

solution Giror voch tot qurope sooroo Data DS zf=sri 3,18, 24, 25, 29, 31, 35, 42,85 orobaro tpp tidano) O first quartiled at= (7.5th value. 7th value + o.so (8th value. 7th value) こ = 35 +0.50 (42-35 ) 03 = 38.5 28_Intes quartile Range! IOR= 63-0 =